OpenStudy (anonymous):

y = a / 1 + be^-x ; (a,b constant) get dy/dx

OpenStudy (bahrom7893):

y = a/(1+be^(-x)) is this what you meant?

OpenStudy (mertsj):

\[y=\frac{a}{1+be ^{-x}}\]

OpenStudy (anonymous):

derivative of \[\frac{1}{f}\] is \[\frac{-f'}{f^2}\]

OpenStudy (anonymous):

\[f(x)=1+b^{-x}\] \[f'(x)=-b^{-x}\ln(b)\]

OpenStudy (anonymous):

should get \[\frac{ab^{-x}\ln(b)}{(1+b^{-x})^2}\]

OpenStudy (anonymous):

oh man was completely wrong. i thought it said \[b^{-x}\] but it says \[be^{-x}\] let me start again

OpenStudy (anonymous):

\[f(x)=1+be^{-x}\] \[f'(x)=-be^{-x}\] derivative is \[\frac{abe^{-x}}{(1+be^{-x})^2}\]